OpenStudy (anonymous):

Which value of x satisfies both 3x + y = 9 and 3x − 5y = 15? 103 23 −1 0

OpenStudy (anonymous):

subtract the equations

OpenStudy (anonymous):

thanks

OpenStudy (kidrah69):

\[ 3x + y = 9 \] \[ -( 3x − 5y = 15)\] We want opposite signs so that the terms will cancel out :) \[3x + y = 9 \] \[ - 3x + 5y = -15)\]

OpenStudy (kidrah69):

Everything cancels out so what would x=?